# Effect of a shear flow on the Darrieus–Landau instability in a Hele-Shaw channel

### v1

@@ 3 Formulation for narrow channels
This is a new result which indicates that curved flames in a Hele-Shaw channel always experience vorticity in the presence of Taylor dispersion, that is whenever $\sqrt{\gamma}\mbox{{Pe}}\neq 0$, even in the absence of viscosity variations as assumed for simplicity in this paper. Equation (10) also implies that the vorticity field remains zero everywhere, except in (the preheat and reaction zones of) the flame where spatial density variations occur. As can be seen from the two-dimensional equations (4)-(7), the Peclet number enters only as $\mbox{{Pe}}^{2}$, a behaviour that was originally pointed out by Zeldovich [21] in a related heat-transfer problem and obtained in other combustion studies such as [22]. Therefore whether the shear flow is in the positive or negative $x$-directions is unimportant in the narrow-channel limit, under the simplifying constant-viscosity assumption adopted. This directional symmetry of the imposed flow is broken when next-order corrections to (4)-(7) are taken into account [20]. The reader may also refer to the recent study [23] on the flame hydrodynamic instabilities based on a different two-dimensional (Euler–Darcy) model taking into account viscosity variations.
@@ 6 Time-dependent numerical simulations
Illustrative results for six cases are presented in Fig. 5. Shown are instantaneous temperature fields and streamlines in the frame moving with the flame adopted at selected times long after the instability has developed.

### v3

@@ 3 Formulation for narrow channels
This is a new result which indicates that curved flames in a Hele-Shaw channel always experience vorticity in the presence of Taylor dispersion, that is whenever $\sqrt{\gamma}\mbox{{Pe}}\neq 0$, even in the absence of viscosity variations as assumed for simplicity in this paper. Equation (10) also implies that the vorticity field remains zero everywhere, except in (the preheat and reaction zones of) the flame where spatial density variations occur. As can be seen from the two-dimensional equations (4)-(7), the Peclet number enters only as $\mbox{{Pe}}^{2}$, a behaviour that was originally pointed out by Zeldovich [21] in a related heat-transfer problem and obtained in other combustion studies such as in [22] in the case of narrow channels. Therefore whether the shear flow is in the positive or negative $x$-directions is unimportant in the narrow-channel limit, under the simplifying constant-viscosity assumption adopted herein. This directional symmetry of the imposed flow is broken when next-order corrections to (4)-(7) are taken into account [20] or also when viscosity variations are taken into account. The reader may also refer to the recent study [23] on the flame hydrodynamic instabilities based on a different two-dimensional (Euler–Darcy) model taking into account viscosity variations.
@@ 6 Time-dependent numerical simulations
Illustrative results for six cases are presented in Fig. 5. Shown are instantaneous temperature fields and streamlines in the frame moving with the flame adopted at selected times long after the instability has developed. The double-headed arrow is in the direction of the imposed shear flow and indicates the equivalence of two opposite parallel directions of the flow. As emphasised at the end of section 3, the later equivalence ceases to hold when the constant-viscosity assumption is relaxed (see, for instance [23]) or in wider channels (see, e.g., [20]).